If Caroline’s investment were expected to yield an annual cash flow of GHS500 for each of the next ten years, assuming discount rate is 5%, (i) Determine the present value of her investment. (ii) Assuming that cash flows were only realised at the beginning of each period, determine the present value of the annuity due. 


1
Expert's answer
2021-04-18T19:23:56-0400

i)


"500\times(1+0.05)^{10}=814"

ii)


"814-500=314"